AAPL wrote:GMAT Prep
A basket contains 5 apples, of which 1 is spoiled and the rest are good. If Herny is to select 2 apples from the basket simultaneously and at random, what is the probability that the 2 apples selected will include the spoiled apple?
A. 1/5
B. 3/10
C. 2/5
D. 1/2
E. 3/5
OA C
We need to determine the probability of selecting a spoiled apple and a non-spoiled apple when selecting two apples.
The number of ways to select the spoiled apple is 1C1 = 1. The number of ways to select one good apple is 4C1 = 4. Thus, the spoiled apple and a good apple can be selected in 1 x 4 = 4 ways.
The number of ways to select 2 apples from 5 is 5C2 = (5 x 4)/2! = 20/2 = 10.
Thus, the probability of selecting the spoiled apple and a good apple is 4/10 = 2/5.
Alternate Solution:
There are two outcomes that satisfy the requirement that the spoiled apple (S) is chosen along with a good apple (G), either S-G or G-S. The probability of S-G is (1/5)(4/4) = 1/5. The probability of G-S is (4/5)(1/4) = 1/5. Since either outcome satisfies our requirement, we add these two probabilities: 1/5 + 1/5 = 2/5.
Answer: C
Scott Woodbury-Stewart
Founder and CEO
[email protected]
See why Target Test Prep is rated 5 out of 5 stars on BEAT the GMAT. Read our reviews

